OpenStudy (anonymous):

Consider the equation: x^2+xy+y^2=1. Find dy/dx Find all points where the tangent line is horizontal.

OpenStudy (anonymous):

I actually found the dy/dx: (-2x-y)/(x+2y) I cannot figure out how to find the horizontal tangents by setting to zero. I'm having a hard time solving.

OpenStudy (freckles):

checking one sec

OpenStudy (freckles):

ok your y' looks great

OpenStudy (freckles):

horizontal lines have slope 0

OpenStudy (freckles):

so we need to find when y'=0

OpenStudy (freckles):

\[\frac{-2x-y}{x+2y}=0\] by the way y' does not exist when x+2y=0 so find when -2x-y=0 and find when x+2y=0 so y'=0 when -2x-y=0 but exclude any numbers from the equation x+2y=0

OpenStudy (freckles):

if you are there can you tell me what you get when you solve -2x-y=0 for y?

OpenStudy (anonymous):

sorry, I was working on some other problems. ok, so I have y= -2x

OpenStudy (freckles):

right and solving the equation in which gives us y' does not exist is x+2y=0 so y=-1/2 x there y=-1/2x and y=-2x share a common point (0,0) right?

OpenStudy (freckles):

so any point of the form (x,-2x) where x is not zero is where you have y' is 0

OpenStudy (freckles):

does that make sense to you?

OpenStudy (freckles):

i must leave ya now but good luck

OpenStudy (freckles):

oh i can wait a sec if you have something else to say

OpenStudy (anonymous):

Thank You! I think I am understanding it now. Thanks!

OpenStudy (freckles):

I just found when the top is zero and excluded any numbers in that set if they were also a number that made the bottom zero

OpenStudy (freckles):

number I meant point

OpenStudy (freckles):

anyways have fun! :)

OpenStudy (freckles):

and actually you might want to see if that is point or see for which points (x,-2x) where x doesn't equal 0 satisfies x^2+xy+y^2=1 x^2+x(-2x)+(-2x)^2=1 x^2-2x^2+4x^2=1 solve for x